OpenStudy (anonymous):

Betty purchased two hotdogs and one soda for $11. Jose and his family purchased two hotdogs and three sodas for $17. Write an equation for Betty’s purchase using “h” for the price of a hotdog and “d” as the price of a soda. Do the same for Jose’s purchase. Determine the cost of a hotdog and a soda.

OpenStudy (anonymous):

Can Someone do the whole problem and show explanation??

OpenStudy (diyadiya):

Betty purchased two hotdogs and one soda for $11 2h+1s=11 Jose and his family purchased two hotdogs and three sodas for $17 2h+3s=17

OpenStudy (anonymous):

2h+d=11 2h+3d=17 ... is this what you're asking ? :P

OpenStudy (hoblos):

for Betty: 2h+d =11 for Jose: 2h+3d =17 can solve this system now?

OpenStudy (diyadiya):

Oops sorry that was d instead of s :D Betty purchased two hotdogs and one soda for $11 2h+1d=11 Jose and his family purchased two hotdogs and three sodas for $17 2h+3d=17

OpenStudy (diyadiya):

Can you solve now?

OpenStudy (anonymous):

can u solve it and show explanation im learning how to solve problems thats why i want to see how the people do it

OpenStudy (diyadiya):

You can use either Substitution or elimination method here Subtract the two equations 2h+3d =17 2h+ d =11 ____________ 0+2d= 6 2d=6 divide both sides by2 2d/2=6/2 d=3

OpenStudy (anonymous):

ok so then what comes up next?

OpenStudy (anonymous):

is that it?is that the answer?

OpenStudy (diyadiya):

NO now we have to find h we have just found out d (that's the price of soda) Now put the value of d in any of the equation and find out h

OpenStudy (diyadiya):

2h+ d =11 2h+ (3) =11 subtract 3 from both sides 2h + 3 -3 = 11 -3 2h=8 divide by 2 2h/2 =8/2 h=8/2 h= ?

OpenStudy (anonymous):

oh ok so we already found the answer for price Of soda now we just need to find h Right?

OpenStudy (diyadiya):

Yup =D

OpenStudy (anonymous):

so d =11?

OpenStudy (anonymous):

h=17 correct?

OpenStudy (diyadiya):

Nope

OpenStudy (diyadiya):

i just solved for d see Subtract the two equations 2h+3d =17 2h+ d =11 ____________ 0+2d= 6 2d=6 divide both sides by2 2d/2=6/2 d=3 so d=3 got it?

OpenStudy (anonymous):

oh yes i go it

OpenStudy (diyadiya):

Now since we have found out d we can find h from any of the above equation by subtituting the value of d so now i take this equation 2h+ d =11 we know the value of d (d=3) So 2h+ 3 =11 Now can you find h?

OpenStudy (anonymous):

i got h=4

OpenStudy (diyadiya):

Yes thats Right =)

OpenStudy (anonymous):

thanx alot for the help but how can i type in my answer for this question?

OpenStudy (diyadiya):

type where?